# Edit Once, Update Everywhere: A Simple Framework for Cross-Lingual Knowledge Synchronization in LLMs

### v1

@@ A.3 Training Data Statistics
Appendix B Implementation Details
We employ EasyEdit Wang et al. (2023c) to implement all the baselines with the default settings. We employ llamafactory Zheng et al. (2024) to implement the Cross-lingual Edition Instruction Tuning (XE-IT) phase of our method. When training on the English editing-only subset, the duration is approximately 10 to 15 hours. Hyperparameters of our X-KDE are in Table 7.
@@ D.1 Detailed Results of Llama2-7b-chat
Table 11: Cross-lingual editing performance of different methods, in terms of F1 score on Qwen2.5-7B-Instruct backbones. Results in green indicates the best results. “Avg.” represents the overall mean of all metrics evaluated across the two languages.

### v2

@@ A.3 Training Data Statistics
Appendix B Experimental Setup
B.1 Alignment Score Calculation
We employ NLLB-600M-distilled [^1] https://huggingface.co/facebook/nllb-200-distilled-600M as the multilingual translation model. And since the translation model is trained on a large-scale parallel dataset in both source and target languages ($X$ and $Y$) by maximizing the conditional generation probability $P(Y|X;\theta)$. Higher alignment between $Y$ and $X$ results in a higher conditional generation probability $P(Y|X)$. To compute the "alignment" score, we input the model’s response and force-decode the ground-truth $Y$ in target language.
B.2 Implementation Details
All experiments were executed on 8 NVIDIA A100 GPUs (80G). We employ EasyEdit wang2023easyedit to implement all the baselines with the default settings. We employ llamafactory zheng2024llamafactory to implement Cross-lingual Edition Instruction Tuning (XE-IT) phase of our method. When training on the English editing only subset, the duration is approximately 10 to 15 hours. Hyper-parameters of our X-KDE are in Table 7.
@@ D.1 Detailed Results of Llama2-7b-chat
Table 11: Cross-lingual editing performance of different methods on Qwen2.5-7B-Instruct backbones. Results in green indicates the best results. “Avg.” represents the overall mean of all metrics evaluated across the two languages.